# Privacy-Preserving Resilient Vector Consensus

### v1

@@ 3.1 Network Model
We consider a multi-agent system with $n$ agents modeled by a time-varying directed graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t)),\text{ }t=0,1,2,\ldots$, where $\mathcal{V}=\{1,2,\ldots,n\}$ is the set of agents, and $\mathcal{E}(t)\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. The set of in-neighbors of an agent $i$ is denoted by $\mathcal{N}_{i}^{\text{in}}(t)=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}(t)\}$, while the set of out-neighbors is $\mathcal{N}_{i}^{\text{out}}(t)=\{j\in\mathcal{V}|(i,j)\in\mathcal{E}(t)\}$, and the corresponding out-degree is $d^{\text{out}}_{i}(t)=|\mathcal{N}_{i}^{\text{out}}(t)|$ with $d^{\text{out}}_{\text{max}}=\max\limits_{\{i,t\}}(d^{\text{out}}_{i}(t))$.
@@
The system has two types of agents, i.e., normal agents and faulty agents. Each normal agent updates its state through local interaction based on predefined rules. Conversely, faulty agents can behave arbitrarily and unpredictably. They are classified into two categories: Byzantine agents and malicious agents. Byzantine agents send different arbitrary states to different neighbors, while malicious agents can only send the same state to all neighbors [11]. Malicious agents are actually a specific type of Byzantine agents. Therefore, for a more general case, we assume that the faulty agents in this paper are Byzantine agents. We denote by $\mathcal{F}\in\mathcal{V}$ the set of faulty agents, and by $f=|\mathcal{F}|$ the total number of faulty agents with $f\leq F$. For any $i\in\mathcal{F}$ and $j\in\mathcal{V}$, we denoted by $\hat{x}_{i}^{j}(t)$ the state $i$th faulty agent sent to $j$th agent at iteration $t$. Let $\overline{\mathcal{V}}=\mathcal{V}-\mathcal{F}$, and $\overline{n}=n-f$, where $\overline{\mathcal{V}}$ is the set and $\overline{n}$ is the number of all normal agents. Without loss of generality, we suppose $\overline{n}$ ones in front are the normal agents, meaning that $\overline{\mathcal{V}}=\{1,2,\ldots,\overline{n}\}$. Besides, we denote by a $\overline{n}\times d$ matrix $\overline{\mathbf{x}}(t)=[x_{1}(t),x_{2}(t),\ldots,x_{\overline{n}}(t)]^{\top}$ the states of all normal agents, where $x_{i}(t)$ denotes the state of a normal agent, and by $\overline{\mathbf{x}}_{0}$ the initial states of normal agents. We define the topology of normal agents by a time-varying directed graph $\overline{\mathcal{G}}(t)=(\overline{\mathcal{V}},\overline{\mathcal{E}}(t))$, where ¯$\overline{\mathcal{E}}(t)\subseteq{\overline{\mathcal{V}}}\times\overline{\mathcal{V}}$. For each normal agent $i\in\overline{\mathcal{V}}$, we denote by $\overline{\mathcal{N}}_{i}^{\text{in}}(t)$ and $n_{f_{i}}(t)$ its set of in-neighbors in $\overline{\mathcal{G}}(t)$ and the number of its faulty in-neighbors in $\mathcal{G}(t)$, respectively.
@@ 3.2 Algorithm Design
We consider a modification of the existing ADRC Using Centerpoint (ADRC) algorithm, i.e., Privacy-Preserving ADRC (PP-ADRC) algorithm. Its details are presented in the Algorithm 1. To protect privacy, zero-mean, decaying Gaussian noise is added to the state of each normal agent during the communication phase. In ADRC, the key to achieving resilience lies in calculating the centerpoint $s_{i}(t)$ during the calculation phase.
@@
  Input: $\mathcal{G}(t)$, $\overline{\mathbf{x}}(t)$, $\Sigma(t)=\lambda^{2}\upsilon^{2t}I_{d}$, $\gamma_{i}(t)$, threshold $\mathrm{thre}$
  Output: final value $x(\infty)$
@@
     Update its state following:
@@
  If $||x_{i}(t+1)-x_{i}(t)||_{2}<\mathrm{thre},\forall i\in\overline{\mathcal{V}}$, the iteration terminates.
For a set of $n$ points in general position in $\mathbb{R}^{d}$, there always exists a centerpoint $p$, which is guaranteed to lie within the convex hull of any subset containing more than $\frac{nd}{d+1}$ points from the given point set [23, 24]. Therefore, if the fraction of faulty agents in a normal agent’s neighborhood is lower than $\frac{1}{d+1}$, the centerpoint always lies in the convex hull formed by the normal agents’ state vectors. However, the fraction $\frac{1}{d+1}$ is just a theoretical value. When the dimensionality exceeds three, calculating a centerpoint becomes a coNP-Complete problem, and we can only use approximation algorithms. Therefore, in practice, if $d>3$, we can only achieve a fraction of $\Omega\left(\frac{1}{d^{2}}\right)$ [17].
It should be pointed out that we choose the noise with zero mean Gaussian distribution with covariance matrix $\Sigma(t)$ for two reasons: 1) The properties of zero-mean and decaying variance are the necessary conditions to reach resilient vector consensus; 2) Gaussian distribution, whose pdf is also proportional to the $2$-norm, facilitates our subsequent privacy analysis when we use the distance between agents (measured by the $2$-norm) as a parameter for evaluating privacy preservation. Additionally, Gaussian noise has a smaller variation in magnituede with dimension changes compared to Laplace noise.
@@ 3.3 Problem of Interests
Different from the resilient vector consensus without adding noise, the randomness of noise makes the result under PP-ADRC become non-deterministic. Thus, we need to characterize metrics for resilient vector consensus and privacy preservation after noise addition, which are provided below.
@@
  • Agreement: For any pair of normal agents $i,j\in\overline{\mathcal{V}}$, there always holds $\lim\limits_{t\rightarrow\infty}\mathbb{E}[||x_{i}(t)-x_{j}(t)||_{2}^{2}]=0$.
To characterize the privacy preservation of the state vector sequences, we first denote by three matrices $\overline{\mathbf{x}}(t),\ \overline{\boldsymbol{\eta}}(t),\ \text{and}\ \overline{\mathbf{y}}(t)\in\mathbb{R}^{\overline{n}\times d}$ the states, noises, and noised states of all the agents at time $t$, respectively. Then, we define the sequences of matrices $\mathbf{X}=\{\overline{\mathbf{x}}(t)\}_{t=0}^{\infty},\mathbf{N}=\{\overline{\boldsymbol{\eta}}(t)\}_{t=0}^{\infty},\ \text{and}\ \mathbf{Y}=\{\overline{\mathbf{y}}(t)\}_{t=0}^{\infty}$ from $t=0$ to $\infty$, and they are in the sample space $\Omega=(\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}}$. Given an initial state $\overline{\mathbf{x}}_{0}$, the sequences $\mathbf{X}$ and $\mathbf{Y}$ are uniquely determined by the noise sequence $\mathbf{N}$ and we define the corresponding function as $\mathbf{Y}_{\overline{\mathbf{x}}_{0}}(\mathbf{\mathbf{N}})=\mathbf{\mathbf{X}}$.
@@
(CGP in Resilient Vector Consensus): Given any pair of initial states $\overline{\mathbf{x}}_{0}$, $\overline{\mathbf{x}}_{0}^{\prime}$, and $\rho>0$, the system is said to satisfy $\rho-$CGP if and only if for all $\alpha>1$
@@
After noise adding, convergence analysis and privacy quantization become challenging. For convergence analysis, we can only describe the situation in terms of expectation. But for the final value, although the expectation is ensured to lie within the convex hull of normal agents’ initial states, we cannot obtain an analytical solution. Specifically, the exact result of each convergence and its deviation from the expectation are unknown. Additionally, characterizing the change of the convex hull after adding noise is very challenging because the vertices and boundaries are constantly changing. Without analytical expressions, it is hard to quantitatively describe the changes in the convex hull. For privacy preservation, we add an infinite amount of noise, and the output is an infinite sequence of state matrices without an explicit distribution expression. Analyzing the Rényi divergence of such infinite matrix sequences is inherently complex and necessitates a specialized approach. Consequently, we are interested in the following questions of the system performing PP-ADRC:
@@
    How can we derive the upper bound of $\rho$-CGP without knowing the specific distribution of two infinite matrix sequences? What are the advantages compared to differential privacy?
@@ 4.1 Resilient Vector Consensus
We first formulate the state evolution of all normal agents as a linear time-varying (LTV) system by leveraging the properties of the centerpoint. Then, by sequentially multiplying the iterative formula of the LTV system from the initial time step, we can leverage repeated reachability and the characteristics of stochastic matrices to establish the resilient vector consensus of PP-ADRC [2].
@@
$\overline{\mathbf{x}}(t+1)=\mathbf{M}(t)\overline{\mathbf{x}}(t)+\mathbf{H}(t)\mathbf{v}(t),\ t=0,1,2,\ldots$ (7)
where $\mathbf{M}(t)\in\mathbb{R}^{\overline{n}\times\overline{n}}$ is a row-stochastic matrix with $[\mathbf{M}(t)]_{ii}=1-\gamma_{i}(t)$, $\mathbf{H}(t)\in\mathbb{R}^{\overline{n}\times\overline{n}}$ has zero diagonal entries ans differs from $\mathbf{M}(t)$ only in the diagonal elements, and $\mathbf{v}(t)=\left[\eta_{1}(t),\eta_{2}(t),\ldots,\eta_{\overline{n}}(t)\right]^{\top}$.
@@
Theorem 1
Under PP-ADRC, if equation (6) holds and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then for any $i\in\overline{\mathcal{V}}$, there exists a random vector $x(\infty)=[x_{i,1}(\infty),\ldots,x_{i,d}(\infty)]^{\top}$ such that $\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i}(t)-x(\infty)||_{2}^{2}\right]=0$, and $\mathbb{E}[x_{i}(t)]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$, where $t=0,1,2,\ldots$.
Proof 4.2.
Here. we use ergodicity333 Ergodicity implies that for the backward product of a row-stochastic matrix sequence, as time approaches infinity, each row becomes identical. to show the convergence. With equation (8), by using the independence and zero-mean of noise, for any $i\in\overline{\mathcal{V}}$, we obtain
@@
Hence, there must exist a random variable $x_{i,k}(\infty)$ satisfying $\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i,k}(t)-x_{i,k}(\infty)||_{2}^{2}\right]=0$, i.e.,
$\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i}(t)-x(\infty)||_{2}^{2}\right]=0,\ \forall i\in\overline{\mathcal{V}}.$ (10)
Additionally, by the zero-mean and independence of noise, we obtain $\mathbb{E}\left[\overline{\mathbf{x}}(t+1)\right]=\mathbb{E}\left[\Psi_{t+1,0}\overline{\mathbf{x}}(0)\right]$. As $\Psi_{t+1,0}$ is row-stochastic, one derives $\mathbb{E}\left[x_{i}(t+1)\right]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$.
@@
It should be pointed out that equation (6) ensures “safety”, while repeated reachability ensures the ergodicity of backward product $\Psi_{t,t_{s}}$, thereby guaranteeing “agreement”. After adding noise, we can only guarantee the “mean square convergence” under expectation. As the goal is to characterize the effect of noise term by expectation, independence, and zero-mean are the crucial factors, rather than the specific distribution of the noise. This means that resilient vector consensus can be achieved for other noise distributions with independence and zero-mean properties.

### v2

@@ 3.1 Network Model
We consider a multi-agent system with $n$ agents modeled by a time-varying directed graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t)),\text{ }t=0,1,2,\ldots$, where $\mathcal{V}=\{1,2,\ldots,n\}$ is the set of agents, and $\mathcal{E}(t)\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. The set of in-neighbors of an agent $i$ is denoted by $\mathcal{N}_{i}^{\text{in}}(t)=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}(t)\}$, while the set of out-neighbors is $\mathcal{N}_{i}^{\text{out}}(t)=\{j\in\mathcal{V}|(i,j)\in\mathcal{E}(t)\}$.
@@
The system has two types of agents, i.e., normal agents and faulty agents. Each normal agent updates its state through local interaction based on predefined rules. Conversely, faulty agents can behave arbitrarily and unpredictably. They are classified into two categories: Byzantine agents and malicious agents. Byzantine agents send different arbitrary states to different neighbors, while malicious agents can only send the same state to all neighbors [14]. Malicious agents are actually a specific type of Byzantine agents. Therefore, for a more general case, we assume that the faulty agents in this paper are Byzantine agents. We denote by $\mathcal{F}\in\mathcal{V}$ the set of faulty agents. Let $\overline{\mathcal{V}}=\mathcal{V}-\mathcal{F}$, and $\overline{n}=n-|\mathcal{F}|$, where $\overline{\mathcal{V}}$ is the set and $\overline{n}$ is the number of all normal agents. Without loss of generality, we suppose $\overline{n}$ ones in front are the normal agents, meaning that $\overline{\mathcal{V}}=\{1,2,\ldots,\overline{n}\}$. Besides, we denote by a $\overline{n}\times d$ matrix $\overline{\mathbf{x}}(t)=[x_{1}(t),x_{2}(t),\ldots,x_{\overline{n}}(t)]^{\top}$ the states of all normal agents, where $x_{i}(t)$ denotes the state of a normal agent $i\in\overline{\mathcal{V}}$, and by $\overline{\mathbf{x}}_{0}$ the initial states of the normal agents. We define the topology of normal agents by a time-varying directed graph $\overline{\mathcal{G}}(t)=(\overline{\mathcal{V}},\overline{\mathcal{E}}(t))$, where ¯$\overline{\mathcal{E}}(t)\subseteq{\overline{\mathcal{V}}}\times\overline{\mathcal{V}}$. For each normal agent $i\in\overline{\mathcal{V}}$, we denote by $\overline{\mathcal{N}}_{i}^{\text{in}}(t)$ and $n_{f_{i}}(t)$ its set of in-neighbors in $\overline{\mathcal{G}}(t)$ and the number of its faulty in-neighbors in $\mathcal{G}(t)$, respectively.
3.2 Privacy Attackers
Regarding privacy preservation, we consider two types of privacy attackers, as referenced in [28].
  1. 1.
    Privacy attackers are within the network. This type of attacker includes two categories: the honest-but-curious attacker and the malicious attacker. The former follows the pre-defined algorithm just like normal agents but attempts to steal other agents’ privacy through the information received from network interactions. The latter refers to the previously mentioned Byzantine agents, who seek to disrupt consensus while simultaneously stealing the privacy of other agents.
  2. 2.
    Privacy attackers are outside the network, i.e., eavesdroppers. This type of attacker is aware of the network topology and can wiretap communication channels to access the information exchanged between agents.
@@ 3.3 Algorithm Design
We consider a modification of the existing ADRC algorithm, i.e., PP-ADRC, as detailed in Algorithm 1. To preserve privacy, zero-mean decaying Gaussian noise is added to the state of each normal agent during the communication phase. In ADRC, the key to achieving resilience lies in calculating the centerpoint $s_{i}(t)$ during the calculation phase.
@@
  Input: $\mathcal{G}(t)$, $\overline{\mathbf{x}}(t)$, $\Sigma(t)=\lambda^{2}\upsilon^{2t}I_{d}$, $\gamma_{i}(t)$, and number of iterations $T$.
  Output: final value $\xi$
@@
        Update its state following
@@
  end for
Here, we choose Gaussian noise for it is canonical under CGP. We can obtain explicit analytical results compared to Laplace noise. In addition, in resilient vector consensus, there are more advantages that the Gaussian mechanism can bring to the system, which are also detailed as follows.
  • Gaussian noise is easier to implement in non-one-dimensional cases compared to Laplace noise in practical systems, which reduces computational complexity.
  • The magnitude of Gaussian noise grows at a slower rate concerning the dimension $d$. Specifically, the scale of Gaussian noise is $\tilde{O}(\sqrt{d})$, whereas Laplace noise scales as $\tilde{O}(d)$. In resilient vector consensus, where the dimension $d$ can be arbitrarily high, Gaussian noise offers greater flexibility in parameter selection.
  • Gaussian noise is an additive noise, meaning that the sum of two Gaussian noises is still a Gaussian noise. In many application scenarios of resilient vector consensus, such as multi-robot rendezvous, geometric data measurement is involved. However, many geometric measurements inherently contain Gaussian noise. The linearity of Gaussian noise makes it easier for the algorithm to achieve the desired privacy preservation parameters.
We introduce decaying Gaussian noise to protect the initial states of normal agents. Privacy leakage occurs not only during the initial transmission but also at each subsequent iteration, though it diminishes over time. Thus, adding noise with the same intensity at every iteration is unnecessary, and injecting noise only to the initial states is also inadequate. A decaying variance better aligns with our privacy preservation requirements while ensuring system convergence. Using a fixed noise variance would prevent normal agents from reaching consensus, causing their state variances to diverge to infinity as time tends to infinity. Therefore, injecting decaying noise is essential for both privacy preservation and consensus. This is also a common approach in privacy-preserving scalar consensus [24, 29, 30].
3.4 Centerpoint Introduction
We provide a brief introduction to the concept and computation of the centerpoint.
For a set of $n$ points in general position in $\mathbb{R}^{d}$, there always exists a centerpoint $p$, which is guaranteed to lie within the relative interior of the convex hull of any subset containing more than $\frac{nd}{d+1}$ points from the given point set [20]. Therefore, if the fraction of faulty agents in a normal agent’s neighborhood is lower than $\frac{1}{d+1}$, the centerpoint always lies in the relative interior of the convex hull formed by the normal agents’ state vectors. However, the fraction $\frac{1}{d+1}$ is just a theoretical value (This fraction corresponds to the depth of the centerpoint.). When the dimensionality exceeds three, calculating a centerpoint becomes a coNP-Complete problem, where we need approximate algorithms. The specific computation method is as follows.
The computation of the centerpoint varies depending on the dimension. In the two-dimensional case, the prune and search algorithm proposed in [31] can find the ideal centerpoint with time complexity of $O(n)$. For three or more dimensions, the exact algorithm for computing the ideal centerpoint was proposed in [32], with time complexity of $O(n^{d-1})$. Chan demonstrated that centerpoint computation can be formulated as an Implicit Linear Programming problem. However, the algorithm’s complexity grows significantly with higher dimensions, restricting its practicality to low-dimensional cases. In high-dimensional settings, approximation algorithms are necessary, categorized into depth approximation and location approximation. The depth approximation algorithm, proposed in [33], has time complexity of $O(n^{\log d})$. This algorithm projects the point set onto a line and uses an $r$-partition, yielding a centerpoint with depth $\Omega\left(\frac{1}{d^{2}}\right)$. For location approximation, an algorithm proposed in [34] provides a polynomial-time solution. It assigns each point a Gaussian perturbation with scale $\varepsilon$ and then uses the Radial Isotropic Transformation to obtain an $\varepsilon$-approximate centerpoint with depth $\Omega\left(\frac{1}{d}\right)$. Specifically, the algorithm guarantees that the point returned is at most $\varepsilon$ away from any half-space defined by the centerpoint with depth at least $\Omega\left(\frac{1}{d}\right)$ in any direction.
@@ 3.5 Problem of Interests
Different from the resilient vector consensus without adding noise, the randomness of noise makes the result under PP-ADRC become nondeterministic. Thus, we need to characterize metrics for resilient vector consensus and privacy preservation after noise addition, which are provided below.
@@
  • Agreement: For any pair of normal agents $i,j\in\overline{\mathcal{V}}$, it always holds that $\lim\limits_{t\rightarrow\infty}\mathbb{E}[||x_{i}(t)-x_{j}(t)||_{2}^{2}]=0$.
To characterize the privacy preservation of the state vector sequences, we first denote by three matrices $\overline{\mathbf{x}}(t),\ \overline{\boldsymbol{\eta}}(t),\ \text{and}\ \overline{\mathbf{y}}(t)\in\mathbb{R}^{\overline{n}\times d}$ the states, noises, and noisy states of all the agents at time $t$, respectively. Then, we define the sequences of matrices $\mathbf{X}=\{\overline{\mathbf{x}}(t)\}_{t=0}^{\infty},\mathbf{N}=\{\overline{\boldsymbol{\eta}}(t)\}_{t=0}^{\infty},\ \text{and}\ \mathbf{Y}=\{\overline{\mathbf{y}}(t)\}_{t=0}^{\infty}$ from $t=0$ to $\infty$, and they are in the sample space $\Omega=(\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}}$. Since the algorithm is specific, given an initial state $\overline{\mathbf{x}}_{0}$, the sequences $\mathbf{X}$ and $\mathbf{Y}$ are uniquely determined by the noise sequence $\mathbf{N}$ and we define the corresponding function as $\mathbf{Y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})=\mathbf{Y}$.
@@
($\rho$-CGP in Resilient Vector Consensus): Given any pair of initial states $\overline{\mathbf{x}}_{0}$, $\overline{\mathbf{x}}_{0}^{\prime}$, and $\rho>0$, the system is said to satisfy $\rho$-CGP if and only if for all $\alpha>1$
@@
Remark 1
Privacy leakage can occur at both the initial iteration and each subsequent iteration. To address this, we focus on the privacy performance of the infinite matrix sequence of noisy states $\mathbf{Y}$. CGP ensures that for any two initial states $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$, the difference in the distributions of the infinite sequences measured by Rényi divergence is bounded by $\alpha\rho\cdot\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2}$. Essentially, an observer cannot distinguish between $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ based on the output sequences.
After noise addition, convergence analysis and privacy quantization become challenging. For convergence analysis, we can only describe the situation in terms of expectation. But for the final value, although the expectation is ensured to lie within the convex hull of the normal agents’ initial states, we cannot obtain an analytical solution. Specifically, the exact result of each convergence and its deviation from the expectation are unknown. In addition, characterizing the change of the convex hull after adding noise is very challenging because the vertices and boundaries are constantly changing. Without analytical expressions, it is hard to quantitatively describe the changes in the convex hull. For privacy preservation, we add an infinite amount of noise, and the output is an infinite sequence of state matrices without an explicit distribution expression. Analyzing the Rényi divergence of such infinite matrix sequences is inherently complex and necessitates a specialized approach. Consequently, we are interested in the following questions of the system performing PP-ADRC:
@@
    How can we derive the upper bound of $\rho$-CGP without knowing the specific distribution of two infinite matrix sequences? What are the advantages compared to DP?
@@ 4.1 Resilient Vector Consensus
We first formulate the state evolution of all normal agents as a linear time-varying (LTV) system using the properties of the centerpoint. Then, by sequentially multiplying the iterative formula of the LTV system from the initial iteration, we can leverage repeated reachability and the characteristics of stochastic matrices to establish the resilient vector consensus of PP-ADRC [2].
@@
$\overline{\mathbf{x}}(t+1)=\mathbf{M}(t)\overline{\mathbf{x}}(t)+\mathbf{H}(t)\mathbf{v}(t),\ t=0,1,2,\ldots,$ (7)
where $\mathbf{M}(t)\in\mathbb{R}^{\overline{n}\times\overline{n}}$ is a row-stochastic matrix with $[\mathbf{M}(t)]_{ii}=1-\gamma_{i}(t)$, $\mathbf{H}(t)\in\mathbb{R}^{\overline{n}\times\overline{n}}$ has zero diagonal entries and differs from $\mathbf{M}(t)$ only in the diagonal elements, and $\mathbf{v}(t)=\left[\eta_{1}(t),\eta_{2}(t),\ldots,\eta_{\overline{n}}(t)\right]^{\top}$.
Proof 4.1.
If the condition in (6) holds for any normal agent $i\in\overline{\mathcal{V}}$ and $t\geq 0$, then by the Proposition V.1 in [2], we can write $s_{i}(t)$ as a convex combination of its normal in-neighbors’ noisy states
$s_{i}(t)=\sum_{j\in\overline{\mathcal{N}}_{i}^{\text{in}}(t)}a_{ij}(t)y_{j}(t)=\sum_{j\in\overline{\mathcal{N}}_{i}^{\text{in}}(t)}a_{ij}(t)[x_{j}(t)+\eta_{j}(t)],$
where $a_{ij}(t)>0$. We can obtain the update equation
$x_{i}(t+1)$ $=\gamma_{i}(t)\sum_{j\in\overline{\mathcal{N}}_{i}^{\text{in}}(t)}a_{ij}(t)[x_{j}(t)+\eta_{j}(t)]+(1-\gamma_{i}(t))x_{i}(t).$
Finally, the system can be represented as
$\overline{\mathbf{x}}(t+1)=\mathbf{M}(t)\overline{\mathbf{x}}(t)+\mathbf{H}(t)\mathbf{v}(t),\ t=0,1,2,\ldots,$
where $\overline{\mathbf{x}}(t)=[x_{1}(t),\ldots,x_{\overline{n}}(t)]^{\top}$ and $\mathbf{v}(t)=\left[\eta_{1}(t),\eta_{2}(t),\ldots,\eta_{\overline{n}}(t)\right]^{\top}$. Matrices $\mathbf{M}(t)$ and $\mathbf{H}(t)$ are
$[\mathbf{M}(t)]_{ij}=\begin{cases}1-\gamma_{i}(t)&\text{if }i=j\\ \gamma_{i}(t)a_{ij}(t)&\text{if }i\neq j\text{ and }j\in\overline{\mathcal{N}}_{i}^{\text{in}}(t)\\ 0&\text{otherwise,}\end{cases}$ (8)
and
$[\mathbf{H}(t)]_{ij}=\begin{cases}\gamma_{i}(t)a_{ij}(t)&\text{if }i\neq j\text{ and }j\in\overline{\mathcal{N}}_{i}^{\text{in}}(t)\\ 0&\text{otherwise.}\end{cases}$ (9)
@@
Theorem 4.2.
Under PP-ADRC, if the condition in (6) holds and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then for any $i\in\overline{\mathcal{V}}$, there exists a random vector $\xi=[\xi_{1},\ldots,\xi_{d}]^{\top}$ such that $\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i}(t)-\xi||_{2}^{2}\right]=0$, and $\mathbb{E}[x_{i}(t)]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$, where $t=0,1,2,\ldots$.
Proof 4.3.
Here, we mainly use strong ergodicity333 Weak ergodicity implies the rows of a backward product of row-stochastic matrices become identical as time approaches infinity, while strong ergodicity additionally guarantees the existence of this limit [35]. to show the convergence. Considering that weak and strong ergodicity are equivalent for the backward product of row-stochastic matrices [36], we only need to prove weak ergodicity.
Following [37], we first establish a positive lower bound for the nonzero entries of $\mathbf{M}(t)$ for $t\geq 0$. The lower bound for diagonal entries is $1-\gamma_{m}$. For non-diagonal entries, each consists of two positive terms: $\gamma_{i}(t)$, which is lower-bounded by $\gamma_{l}$, and $a_{ij}(t)$, which lacks a direct lower bound. However, under the condition in (6), the centerpoint can be expressed as a convex combination of normal agents’ noisy states, allowing us to determine specific values of $a_{ij}(t)$. Defining $a_{l}$ as the minimum value among all $a_{ij}(t)$, we obtain a lower bound of $\gamma_{l}a_{l}$ for the non-diagonal entries. Thus, the overall lower bound for the nonzero entries of $\mathbf{M}(t)$ is $\min(1-\gamma_{m},\gamma_{l}a_{l})$. Then, by using repeatedly reachable graph sequence, we can establish the weak ergodicity of $\{\mathbf{M}(t),t\geq 0\}$ according to the Theorem V.2 in [2], which leads to the strong ergodicity. Using equation (10), along with the independence and zero-mean properties of the noise and the strong ergodicity, for any $i\in\overline{\mathcal{V}}$ and $k=1,2,\dots,d$, we obtain that
@@
Hence, there must exists a random variable $\xi_{k}$ satisfying $\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i,k}(t)-\xi_{k}||_{2}^{2}\right]=0$, i.e.,
$\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i}(t)-{\xi}||_{2}^{2}\right]=0,\ \forall i\in\overline{\mathcal{V}}.$ (12)
Additionally, by the independence and zero-mean properties of noise, we obtain $\mathbb{E}\left[\overline{\mathbf{x}}(t+1)\right]=\mathbb{E}\left[\Psi_{t+1,0}\overline{\mathbf{x}}(0)\right]$. As $\Psi_{t+1,0}$ is row-stochastic, one derives $\mathbb{E}\left[x_{i}(t+1)\right]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$.
@@
It should be pointed out that the condition in (6) ensures “safety”, while repeated reachability ensures the ergodicity of backward product $\Psi_{t,t_{s}}$, thereby guaranteeing “agreement”. After adding noise, we can only guarantee the “mean square convergence” under expectation. As the goal is to characterize the effect of noise term by expectation, independence, and zero mean are the crucial factors, rather than the specific distribution of the noise. This means that resilient vector consensus can be achieved for other noise distributions with independence and zero-mean properties.
Remark 4.5.
Theorem 4.2 provides only sufficient conditions, making it essential to explore the necessary conditions as well. The key challenge lies in finding the necessary conditions of establishing ergodicity of $\{\mathbf{M}(t),t\geq 0\}$. This requires that the product of $\mathbf{M}(t)$ over a finite iterations has an ergodic coefficient upper bounded by a positive value less than $1$. Given that the topology $\overline{\mathcal{G}}(t)$ is directed and time-varying, and considering that the elements of $\mathbf{M}(t)$ lack analytical expressions due to the centerpoint-based algorithm, existing ergodicity results are inapplicable. This presents a highly challenging problem. Additionally, while [37] provides a uniform lower bound for the nonzero entries of the system matrix to analyze convergence rate, we can only prove the existence of such a bound without deriving an explicit expression. The main challenge lies in analyzing the centerpoint, which is known to lie within the relative interior of the convex hull of the normal agents’ states [2, 20], but its exact location and analytical expression remain unknown. This issue also appears in other resilient vector consensus studies [2, 20, 18], where the absence of an explicit lower bound precludes convergence rate analysis. In our future work, we aim to derive the minimal topological conditions necessary to ensure ergodicity and derive a precise lower bound for the nonzero entries of $\mathbf{M}(t)$.